OpenStudy (anonymous):

Given that f(x) = 5x − 10 and g(x) = x + 3, solve for f(g(x)) when x = −1.

OpenStudy (anonymous):

@Destinymasha destiny can you help him?

OpenStudy (anonymous):

uhh ill try ohh and @Destinymasha any chance ur good in science ?!?!??! or nah

OpenStudy (destinymasha):

SCIENCE:} yes! Math? I sorry - me no good at math

OpenStudy (anonymous):

@Destinymasha I LOVE YOU ONCE AGAIN PLEASE HELP ME IN SCIENCE !!

OpenStudy (destinymasha):

alrighty .......

OpenStudy (anonymous):

ill make question

OpenStudy (destinymasha):

~post it in science, though

OpenStudy (anonymous):

f(x) = 5x - 10 g(x) = x + 3 for x=-1 g(-1)= -1 + 3 = 2 f(g(x)) = 5(2) - 10 = 10 - 10 = 0